Calculate the number of mass of phosphoric acid that is produced when 1.45 x1024 molecules of sulfuric acid react with calcium phosphate.

Ca3(PO4)2 + 3H2SO4 à 3CaSO4 + 2H3PO4
1
Expert's answer
2020-10-14T05:46:09-0400

1 mole of matter contains 6.02214076×10²³ particles.

Given number of molecules represents: 1.45 x1024 / 6.02214076×10²³ = 2.4 moles

n (H3PO4) = 2/3 n (H2SO4) = 2/3 x 2.4 = 1.6 moles

n = m/M

M (H3PO4) = 98 g/mol

m (H3PO4) = n x M = 1.6 x 98 = 156.8 g